# Riemannian Metric Learning: Closer to You than You Imagine

### v1

@@ Acknowledgments
All of these methods require sufficient insight into the structure of the problem, either in terms of the objective or in terms of the parametrization of the metric. This is the reason why they have been applied mostly in approaches considering constant metrics or kernel methods. The general theory of Riemannian metric learning is thus far not developed enough to produce structured methods regardless of the chosen parametrization which can also applied, e.g., to neural network-based approaches, see also Section 5.
@@
  • Statistical and Optimization Guarantees for Nonlinear Metric Learning. Developing rigorous statistical and optimization frameworks to provide theoretical guarantees and performance bounds for Riemannian metric learning methods.
@@
Theoretical advancements in the field of metric learning remain relatively limited compared to the volume of empirical studies in the field. However, there is a notable recent trend toward theoretical and statistical investigations, particularly in the context of preference systems and metric learning [18, 184, 106, 129, 33, 190, 36].
@@
More recently, Canal et al. [33], Wang et al. [190], Chen et al. [36] have explored the intersection of Mahalanobis metric learning and user preference modeling. In these studies, the objective is to jointly learn a Mahalanobis metric and user embeddings $(u_{j})$ in the space, where a user’s preference for $x_{1}$ over $x_{2}$ is encoded as $||x_{1}-u||<||x_{2}-u||$. Initial theoretical results were presented in [33] and later refined by [190], which introduced low-rank assumptions to improve the learning process. Furthermore, [36] extended these ideas to pluralistic alignment in fine-tuning large language models (LLMs), further motivating theoretical advancements in the field.
@@
In the same way, some theory related to the optimization problem exists, but mainly for Mahanalobis distance metric learning [203, 19, 185], scarcely for Riemannian metric learning problems concerning the geodesic distance [30, 148, 23] and nearly nothing for the other Riemannian objects $(\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$ [121, Chapter 4-5]. Theoretical results on the optimization related to these objects would greatly enhance the understanding of the topic.
@@
Numerous Riemannian optimization methods [153, 107] have been developed to account for the geometric structure of the space associated with the learnable parameters. Incorporating geometry into optimization can be compared to employing a Newton method rather than a first-order gradient descent as it leverages additional information about the underlying manifold.
@@
This gap is understandable, given the wide diversity of use cases in Riemannian metric learning, which complicates the design of general-purpose algorithms. Nonetheless, as outlined in Section 3, the associated optimization problem is inherently general. To be more specific, a library specialized in the computation of the Riemannian objects according to the metric parametrization $\mathcal{F}(\mathfrak{g})=(\mathbf{d}_{\mathfrak{g}},\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$ is to be waited. It is reasonable to anticipate that future research will focus on making these methods scalable (Section 4.1.3) akin to advancements observed in deep learning.
@@
All in all, this paper has highlighted the richness and potential of the Riemannian metric learning framework across various applications (Section 3.2). As we have seen, distance metric learning has helped to increase noticeably the performance of classification methods in the past. We begin to see similar effects of Riemannian metric learning in a range of applications, improving the performance in, e.g., generative modeling, trajectory inference, and causal inference. Since we have argued that tools descending from a Riemannian metric are applicable in many contexts in a flexible manner, we anticipate this list of successful applications will grow further as more researchers discover the possibilities of Riemannian metric learning. Thus, we hope that his review provides them with an accessible and motivating introduction and we encourage readers to explore new ways to leverage it in their respective fields.
Despite its promise, Riemannian metric learning still faces significant computational and theoretical challenges (Section 5). While there is substantial theoretical knowledge about the infinite-dimensional Riemannian manifold of Riemannian metrics, it is not connected to any statistical theory or similar for the metric learning problem. From a methodological perspective, research in distance metric learning is already well-established, providing a solid foundation upon which Riemannian metric learning could build (Sections 3.2.1 and 4.2.1). Providing a solid theoretical framework along with flexible computational tools will undoubtedly be demanding, but also offers exciting opportunities to further the adoption and impact.
In conclusion, we firmly believe that Riemannian metric learning represents an exciting new frontier in representation learning—an emerging Eldorado with vast, unexplored potential in both theory and practice.
@@ Appendix A Some Rigorous Aspects of Riemannian Geometry
Typically one assumes that the differentiable structure in Appendix A is maximal with respect to the given conditions. This is can be achieved by extending the structure by parametrizations compatible with condition (2).
@@
One can obtain a coordinate-free formulation, which is also useful for infinite dimensional manifolds, by defining a bilinear operator through $\Gamma(\partial_{i}\psi,\partial_{j}\psi)=\sum_{k=1}^{d}\Gamma^{k}_{ij}\partial_{k}\psi$ and realizing that because of Section 2 this is the unique bilinear form fulfilling the property of the following definition.
[…]

### v2

@@ Acknowledgments
All of these methods require sufficient insight into the structure of the problem, either in terms of the objective or in terms of the parametrization of the metric. This is the reason why they have been applied mostly in approaches considering constant metrics or kernel methods. The general theory of Riemannian metric learning is thus far not developed enough to produce structured methods regardless of the chosen parametrization which can also applied, e.g., to neural network-based approaches, see also Section˜5.
@@
  • Statistical Guarantees for Nonlinear Metric Learning. Developing rigorous statistical and optimization frameworks to provide theoretical guarantees and performance bounds for Riemannian metric learning methods.
@@
Theoretical advancements in the field of metric learning remain limited compared to the volume of empirical studies in the field. However, there is a notable recent trend toward theoretical and statistical investigations, particularly in the context of preference systems and metric learning [20, 191, 109, 132, 36, 197, 39].
@@
[…]
@@
In the same way, some theory related to the optimization problem exists, but mainly for Mahanalobis distance metric learning [210, 21, 192], scarcely for Riemannian metric learning problems concerning the geodesic distance [33, 152, 25] and nearly nothing for the other Riemannian objects $(\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$ [124, Chapter 4-5]. Theoretical results on the optimization related to these objects would greatly enhance the understanding of the topic.
@@
Many Riemannian optimization methods [158, 110] have been developed to account for the geometric structure of the space associated with the learnable parameters. Incorporating geometry into optimization can be compared to employing a Newton method rather than a first-order gradient descent as it leverages additional information about the underlying manifold.
@@
This gap is understandable, given the wide diversity of use cases in Riemannian metric learning, which complicates the design of general-purpose algorithms. Nonetheless, as outlined in Section˜3, the associated optimization problem is inherently general. To be more specific, a library specialized in the computation of the Riemannian objects according to the metric parametrization $\mathcal{F}({\mathfrak{g}})=(\mathbf{d}_{\mathfrak{g}},\mathrm{Exp}_{\mathfrak{g}},\mathfrak{T}_{\mathfrak{g}},\text{vol}_{\mathfrak{g}})$ is to be waited. It is reasonable to anticipate that future research will focus on making these methods scalable (Section˜4.1.3) akin to advancements observed in deep learning.
@@
All in all, this paper has highlighted the richness and potential of the Riemannian metric learning framework across various applications (Section˜3.2). As we have seen, distance metric learning has helped to increase noticeably the performance of classification methods in the past. We begin to see similar effects of Riemannian metric learning in a range of applications, improving the performance in, e.g., generative modeling, trajectory inference, and causal inference. Since we have argued that tools descending from a Riemannian metric are applicable in many contexts in a flexible manner, we anticipate this list of successful applications will grow further as more researchers discover the possibilities of Riemannian metric learning. Thus, we hope that his review provides them with an accessible and motivating introduction and we encourage readers to explore new ways to leverage it in their respective fields.
Despite its promise, Riemannian metric learning still faces significant computational and theoretical challenges (Section˜5). While there is substantial theoretical knowledge about the infinite-dimensional Riemannian manifold of Riemannian metrics, it is not connected to any statistical theory or similar for the metric learning problem. From a methodological perspective, research in distance metric learning is already well-established, providing a solid foundation upon which Riemannian metric learning could build (Sections˜3.2.1 and 4.2.1). Providing a solid theoretical framework along with flexible computational tools will undoubtedly be demanding, but also offers exciting opportunities to further the adoption and impact.
In conclusion, we firmly believe that Riemannian metric learning represents an exciting new frontier in representation learning, with substantial potential for both theoretical advancements and practical applications. However, realizing this potential requires addressing a number of computational and theoretical challenges to ensure scalable methods and rigorous guarantees.
@@ Appendix A Some Rigorous Aspects of Riemannian Geometry
Typically one assumes that the differentiable structure in Appendix˜A is maximal with respect to the given conditions. This is can be achieved by extending the structure by parametrizations compatible with condition (2).
@@
One can obtain a coordinate-free formulation, which is also useful for infinite dimensional manifolds, by defining a bilinear operator through $\Gamma(\partial_{i}\psi,\partial_{j}\psi)=\sum_{k=1}^{d}\Gamma^{k}_{ij}\partial_{k}\psi$ and realizing that because of Section˜2 this is the unique bilinear form fulfilling the property of the following definition.
Definition 19 (Christoffel operator).